Consider the possible merger between Company V and Company Z. The market shares of the companies in the industry are as follows:

Company                   Market Share

Company U                     35%

Company V                      25%

Company W                    20%

Company X                       10%

Company Y                        5%

Company Z                        5%

A. What is the Herfindahl-Hirschman index for this industry before the merger of V and Z?

B. What is the expected change in the HHI if V and Z merge?

C. Would this be considered a merger that would reduce competition? Explain.
Answer #1

1.
HHI is sum of squares of all market shares
Hence, HHI=35^2+25^2+20^2+10^2+5^2+5^2=2400

2.
Expected change=35^2+30^2+20^2+10^2+5^2-2400=250

3.
Yes as the resulting index is more than the threshold of 2500